OpenStudy (anonymous):

If a prism has the triangular base shown and a height of 3", then its volume is 384 cubic inches. true or false....*drawing included*

OpenStudy (anonymous):



OpenStudy (anonymous):

area of the triangle = 0.5 * base * height = 0.5 * 16 * 8 = 64 vol = this * ht = 64*3 = 192

OpenStudy (anonymous):

so its false